# MARM: Unlocking the Future of Recommendation Systems through Memory Augmentation and Scalable Complexity

### v2

@@ 2.1.3. Multi-Layer Target-Attention
6:     $[\mathbf{Iid1},\dots,\mathbf{Iidn}]$ = ModelSparseFeatureLookUp([Iid1, …, Iidn])
@@
[…]
@@
17:     $[\texttt{IidY}^{1}_{\texttt{UidX}},\dots,\texttt{IidY}^{L}_{\texttt{UidX}}]=\texttt{UserItemDepthHash}(\texttt{UidX},\texttt{IidY},[1,\dots,L])$
@@
Up to now, we have obtained the learnable target-attention ID embedding and subsequent un-learnable target-attention cached results. According to the learnable embedding, we first calculate them with respect to a certain target item information $\mathbf{IidY}$ through target attention mechanism:
(5) $\small\begin{split}\mathbf{IidY}_{\texttt{UidX}}^{1}&=\texttt{Target-Attention}(\mathbf{IidY},[\mathbf{Iid1},\dots]),\\ \end{split}$
Next, we feed the results $\mathbf{IidY}_{\texttt{UidX}}^{1}\in\mathbb{R}^{d}$ to calculate with un-learnable cached results to simulate the high-FLOPS masked-self-attention ($\mathcal{O}(n^{2}*d)$) by the low-FLOPS target-attention ($\mathcal{O}(n*d)$):
(6) $\small\begin{split}\mathbf{IidY}_{\texttt{UidX}}^{2}=&\texttt{Target-Attention}^{1}(\mathbf{IidY}_{\texttt{UidX}}^{1},[\mathbf{Iid1}_{\texttt{UidX}}^{1},\dots]),\\ \mathbf{IidY}_{\texttt{UidX}}^{3}=&\texttt{Target-Attention}^{2}(\mathbf{IidY}_{\texttt{UidX}}^{2},[\mathbf{Iid1}_{\texttt{UidX}}^{2},\dots]),\\ &\dots\dots\dots\dots\\ \mathbf{IidY}_{\texttt{UidX}}^{L+1}=&\texttt{Target-Attention}^{L}(\mathbf{IidY}_{\texttt{UidX}}^{L},[\mathbf{Iid1}_{\texttt{UidX}}^{L},\dots]),\\ \end{split}$
where the $\mathbf{IidY}_{\texttt{UidX}}^{L+1}\in\mathbb{R}^{d}$ denotes the final sequences interests modeling output belongs to (UidX, IidY) pair, which will feed to latter multi-task learning module to fit a real label.
2.1.4. Sending results to Cache Memory
[…]
Afterwards, we can send the values with their keys in order:
In streaming training, as we continuously save the real-time computation results, our MARM cache will gradually accumulate the computation results of each layer of the model for all of the user’s historical behaviors. Through the natural accumulation of cache, our MARM successfully applies decoder-only Transformer architecture with minor inference complexity FLOPs cost. The pseudo-code of our MARM is shown in Algorithm 1, and Figure 3 also provides a user-watching example in streaming setting.
2.2. MARM with SIM
In Section 2.1, we have discussed a naive MARM setting, modeling the users’ interests under the latest watched items sequence. However, since the sequence length $n$ in Eq.(1) is limited (about 100), thus the earlier stored results will no longer be used once out of range, resulting in a waste of storage resources. This section further extends MARM with Search-based interests model (SIM)(Pi et al., 2020), to maximize Cached results effectiveness.
[…]
(9) $\footnotesize\begin{split}[\texttt{Sid1}^{i},\dots,\texttt{SidK}^{i}]&=\texttt{MARMGSU}(\texttt{IidY},[\texttt{Iid1},\dots,\texttt{Iid}10000],K,i),\\ \end{split}$
[…]
$\footnotesize\begin{bmatrix}\texttt{Sid1}&\texttt{Sid2}&\texttt{Sid3}&\dots&\texttt{SidK}\\ \texttt{Sid1}^{1}&\texttt{Sid2}^{1}&\texttt{Sid3}^{1}&\dots&\texttt{SidK}^{1}\\ \texttt{Sid1}^{2}&\texttt{Sid2}^{2}&\texttt{Sid3}^{2}&\dots&\texttt{SidK}^{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \texttt{Sid1}^{L}&\texttt{Sid2}^{L}&\texttt{Sid3}^{L}&\dots&\texttt{SidK}^{L}\end{bmatrix}$
where the different layer searched sequence maybe different e.g., $\texttt{Sid}1^{i}\neq\texttt{Sid}1^{i+1}$. Afterwards, we can look up their corresponding cached results to support our MARM module in Figure 4(b).
2.3. MARM to Support Other Models
Actually, the industrial RecSys always introduces several models to filter items candidates layer by layer:
  • Retrieval model: without item candidate, only utilizing user’s latest interaction items to find a thousands items group from the full billion item pool.
  • Cascading model: a small ranking model, utilizing extremely efficient modules to select top hundreds item candidates from the retrieval model generated thousands of items candidates.
  • Ranking model: in a broad sense, it is the most complex model, to find the best dozens items from hundreds of candidates.
Generally, these models are trained using different strategies and are isolated from each other. Therefore, here is an interesting assumption: can such our ranking model MARM cache storage contribute to other Retrieval/Cascading models? Unsurprisingly, we found our MARM results could support them seamlessly:
  • For Retrieval model, since there is no candidate set of items, we use user’s latest 200 watched items to look up the cache results from MARM, and then apply the user ID feature to replace item candidates to conduct target-attention as shown in Figure 4(c).
  • For Cascading model, we implement a same multi-layer target-attention process as our Ranking model in Eq.(6). It is important to note that in the cascading model, we do not perform re-accumulation of the MARM cache; instead, we directly use the already accumulated MARM cache and only used latest short-term sequences of users. This approach saves resources while also enhancing the consistency between cascading and ranking models.
@@ 3.1.1. Datasets
To evaluate the effectiveness of our MARM, we conduct detailed experiments at Kuaishou’s international business (Kwai). To be specific, this short-video scenario includes approximately 30 million users and 62 million short videos, with each user reading an average of 133 short videos per day. We collect past 6 months logs of this scenario, and then conduct the common multi-task learning with naive MARM setting (without SIM) in streaming training manner.
@@ 3.1.2. Metrics
For brevity, we present the GAUC and training loss of the long-view prediction task only to reflect the model’s performance. Indeed, GAUC has been validated on Kuaishou as a metric most closely related to online effectiveness, which employs a weighted average of the AUC for each user, where the weights are determined by the number of samples for that user:
@@ 3.2.2. Scaling up the depth $L$ and sequence length $n$
As shown in Figure 5(b), we can see a significant positive correlation to increase MARM depth $L$ and sequence length $n$. Additionally, it is observed that when the length of the user’s behavior sequence increases, even reaching 6400, increasing the MARM depth $L$ still could lead to a notable improvement in model performance, which validates caching the intermediate results is a potential way to unlock new direction to build better RecSys.
@@ 3.2.3. Scaling up the cache size $\mathcal{C}$
Based on the above studies, we further analyzed the relationship between the model’s performance with the cache size C. Specifically, we draw several lines to represent same Cache size and FLOPs variants but with different settings, e.g., 400*128 has three variants 4*100*128, 2*200*128, and 1*400*128. As shown in Figure 5(c), there is a noticeable power-law improvement trend in model performance when the cache size is increased. An interesting phenomenon is that when the cache size is small, increasing the sequence length $n$ yields significantly better results than increasing the depth $L$, e.g., 200*128. However, once the cache size reaches a certain level, the effects of increasing sequence length and MARM depth become comparable, which leads to an excited observation that models with same enough Cache size will show similar performance, e.g., 6400*128.
@@ 3.3. Real-World Experiments
This section focuses on answering the following research questions:
  • RQ1: How does MARM perform compared to other sequence modeling algorithms?
  • RQ2: How does MARM bring improvements to the industry ensemble model?
  • RQ3: How does the SIM idea work as expected in MARM?
  • RQ4: What are the costs associated with MARM?
  • RQ5: How does MARM contribute to the online gains?
@@
TWIN $0.8249\pm 0.00005$ $0.7142\pm 0.00009$ 500M
TWIN V2 $0.8251\pm 0.00023$ $0.7147\pm 0.00007$ 830M
MARM(L=1) $0.8256\pm 0.00008$ $0.7157\pm 0.00006$ 700M $\leftarrow$ 10.4B*
MARM(L=2) $0.8262\pm 0.0004$ $0.7169\pm 0.00012$ 900M $\leftarrow$ 20.6B*
MARM(L=4) $\mathbf{0.8270\pm 0.0003}$ $\mathbf{0.7183\pm 0.00008}$ 1.3B $\leftarrow$ 40.8B*
@@ 3.3.1. Compared Approaches Details
Generally speaking, MARM can be seen as a user sequence modeling module, therefore we select the following strong methods to verify MARM ability:
  • Baseline: A model using the multi-task MoE structure with user features, video features, and statistical features, where the user sequence information is the result of sum pooling of the user’s short-term historical behavior.
@@
  • SIM Soft(Pi et al., 2020): A two-stage modeling approach with GSU-ESU, where GSU uses pre-trained multi-modal embeddings of videos to calculate the inner product and select the top-k most relevant videos from the user’s history. The user history length we are using here is 15000.
  • TWIN(Chang et al., 2023): A two-stage modeling approach under the SIM architecture, aligning the computation methods of GSU and ESU to enhance consistency between the two stages. The user history length we are using is 15,000.
  • TWIN V2(Si et al., 2024): This approach uses hierarchical clustering to reduce the scale of ultra-long user historical lengths, followed by modeling using the TWIN method. The user history length we are using is 100,000.
  • HSTU*(Zhai et al., 2024): This method employs an HSTU-style multi-layer self-masked attention mechanism to model user history (time complexity $\mathcal{O}(n^{2}*d)$). Due to computational limitations, each historical item only uses a few features: item ID, author ID, tag, and user feedback, with 2000 history length and 4 depth. Its FLOPs can be seen as the un-cache version MARM.
@@
  • Ensemble experiment: After a modification is verified to be effective, such method will be fused with latter modifications.
3.3.2. Individual Experiment Performance Comparisons (RQ1)
@@
      +TWIN $0.8260\pm 0.00005$ $0.7156\pm 0.00006$ 508M
         +TWIN V2 $0.8268\pm 0.00023$ $0.7167\pm 0.00011$ 838M
            +HSTU* $0.8268\pm 0.00017$ $0.7168\pm 0.00021$ 5.5B
@@
Table 3. Online A/B testing results of Short-Video services at Kwai.
@@
3.3.3. Ensemble Experiment Offline Performance (RQ2)
@@
This section aims to verify the following question: could MARM module provide further improvements to a model that already incorporates various long-term and short-term user behavior modeling? Therefore, we conducted another set of ensemble experiments and the results are shown in Table 2. Specifically, we sequentially added the DIN, SIM Soft, TWIN, TWIN V2, and MARM modules to the baseline, observing the improvements brought by each new module. From this Table, after adding a series of long-term and short-term modeling modules, adding a simplified HSTU-style module did not lead to significant improvements, possibly because the features we used, attention depth, or historical length did not reach a certain scale. Besides, another reason might be that our MARM could save more interaction sample knowledge. While a pure item sequence has only fewer tens of attributes, the MARM’s cached results are always thousand-dimension compressed target item query information, which achieves potential sample communication within MARM. In addition, directly adding an HSTU-style module significantly increased the computational burden on our model. In our scenario, when we add MARM to a system that includes several long-term and short-term sequence modeling models, there is still a significant improvement in accuracy. MARM is a very practical module that can be added to most recommendation models at a very controllable cost.
3.3.4. GSU searched Top-K Overlapped Rate of MARM (RQ3)
@@
This section aims to understand how the attention depth of MARM can lead to sustained improvements. In the framework of MARM with SIM, each layer of the MARM Block models the GSU and ESU phases in a manner similar to TWIN. The GSU returns the top-k user histories that are most relevant to the target item. As shown in Figure 6, we analyzed the overlap of the histories returned by the GSU of the four-layer MARM Block and found that the proportion of independence for each layer is relatively high, exceeding 50%. Moreover, the direct overlap between any two layers is less than 20%. This indicates that each layer of MARM focuses on different historical content, forming a high-level expression of interests.
3.3.5. MARM Cost Discussion (RQ4)
@@
As an equivalent form of an un-cached MARM, HSTU follows a first sample-aggregation then learning paradigm to alleviate calculation pressure, its overall computational complexity for a user with a lifelong historical behavior is about $O(L*n^{2}*d*(\frac{N}{r}))$, where N denote the total interaction amount of a user, and $\textit{r}<<n$ is the aggregation level of user samples, meaning that every r user samples are aggregated together for training. In practice, a larger aggregation r typically reduces the modeling complexity but can lead to delays in user feedback entering the model training, and easily results in a decline in online metrics, creating a trade-off between cost and effectiveness.
@@
Compared to the above HSTU applies a high-FLOPs masked-self-attention, our MARM is a more low-resource solution, which does not need to aggregate training samples for user modeling. For the overall learning process of a user, the computational complexity of the MARM is about $O(L*n*d*N))$. As a replacement, MARM does require additional storage, which we previously referred to as cache size $\mathcal{C}$. In our scenario, with an attention depth $L=4$ and a sequence length $n=6000$, the storage used by MARM is 60TB. The combined storage and increased computational overhead is approximately $1/8$ that of a direct multi-layer self-attention approach.
3.3.6. Online Results (RQ5)
@@
To quantify the contribution of the MARM model to Kuaishou’s international business (Kwai), we implemented it across the Retrieval, Cascading, and Ranking stages, validating its effectiveness through our online A/B testing system. The model’s performance was assessed based on key metrics, including core playtime metrics and interaction metrics such as average watch time per use and the number of likes. Table 3 presents the online results of MARM across various recommendation stages.
@@
Our app’s retention metrics are positive, indicating that the launch of MARM has resulted in a better user experience.Notably, during multiple stage, MARM achieved a significant increase in average user watch time of +2.079%, highlighting its crucial role in enhancing user watch time. While there were slight negative impacts on Comment/Forward/Follow, significant gains were achieved in user interactions through Like. Since the primary goal of the recommendation system is to maximize average watch time, the online results still within a reasonable range of substitution.

### v3

@@ 2.3. MARM to Support Other Models
6:   $\mathbf{Iid1},\dots,\mathbf{Iidn}]$ = ModelSparseFeatureLookUp([Iid1, …, Iidn])
@@
8:    $\texttt{Iid1}^{i}_{\texttt{UidX}},\dots,\texttt{Iidn}^{i}_{\texttt{UidX}}]$ = UserItemDepthHash(UidX, [Iid1, …], ii)
9:    $\mathbf{Iid1}^{i}_{\texttt{UidX}},\dots,\mathbf{Iidn}^{i}_{\texttt{UidX}}]$ = MARMCacheLookUp($\texttt{Iid1}^{i}_{\texttt{UidX}},\dots]$)
@@
17:   $\texttt{IidY}^{1}_{\texttt{UidX}},\dots,\texttt{IidY}^{L}_{\texttt{UidX}}]=\texttt{UserItemDepthHash}(\texttt{UidX},\texttt{IidY},[1,\dots,L])$
@@ 3.1.1. Datasets
To evaluate the effectiveness of our MARM, we conduct detailed experiments at Real-world RecSys. To be specific, this short-video scenario includes approximately 30 million users and 62 million short videos, with each user reading an average of 133 short videos per day. We collect past 6 months’ logs of this scenario, and then conduct the common multi-task learning with naive MARM setting (without SIM) in streaming training manner.
@@ 3.1.2. Metrics
For brevity, we present GAUC and training loss of long-view prediction task only to reflect model’s performance. GAUC, a metric most closely related to online effectiveness, employs a weighted average of the AUC for each user, where the weights are determined by the number of samples for that user:
@@ 3.2.2. Scaling up the depth $L$ and sequence length $n$
As shown in Figure 5(b), we can see a significant positive correlation between increasing MARM depth $L$ and sequence length $n$. Additionally, it is observed that when the length of the user’s behavior sequence increases, even reaching 6400, increasing the MARM depth $L$ still could lead to a notable improvement in model performance, which validates that caching the intermediate results is a potential way to unlock a new direction to build better RecSys.
@@ 3.2.3. Scaling up the cache size $\mathcal{C}$
Based on the above studies, we further analyzed the relationship between the model’s performance with the cache size C. Specifically, we draw several lines to represent same Cache size and FLOPs variants but with different settings, e.g., 400*128 has three variants 4*100*128 $\leftrightarrow$ 2*200*128 $\leftrightarrow$ 1*400*128. As shown in Figure 5(c), there is a noticeable power-law improvement trend in model performance when the cache size is increased. An interesting phenomenon is that when the cache size is small, increasing the sequence length $n$ yields significantly better results than increasing the depth $L$, e.g., 200*128. However, once the cache size reaches a certain level, the effects of increasing sequence length and MARM depth become comparable, which leads to an excited observation that models with same enough Cache size will show similar performance, e.g., 6400*128.
@@ 3.3. Real-World Experiments
TWIN $0.8249\pm 0.00005$ $0.7142\pm 0.00009$ 90M
TWIN V2 $0.8251\pm 0.00023$ $0.7147\pm 0.00007$ 150M
MARM(L=1) $0.8256\pm 0.00008$ $0.7157\pm 0.00006$ 127M $\leftarrow$ 2.7B*
MARM(L=2) $0.8262\pm 0.0004$ $0.7169\pm 0.00012$ 164M $\leftarrow$ 5.4B*
MARM(L=4) $\mathbf{0.8270\pm 0.0003}$ $\mathbf{0.7183\pm 0.00008}$ 238M $\leftarrow$ 10.5B*
@@ 3.3.1. Compared Approaches Details
Generally speaking, MARM can be seen as a user sequence modeling module; therefore, we select the following strong methods to verify MARM’s ability:
  • Baseline: A model using multi-task MoE structure with features of users, videos, and statistics. User sequence information is the result of sum pooling of user’s short-term historical behavior.
@@
  • SIM Soft(Pi et al., 2020): A two-stage modeling approach with GSU-ESU, where GSU uses pre-trained multi-modal embeddings of videos to calculate inner product and select top-k most relevant videos from user’s history. The user history length used here is 15000.
  • TWIN(Chang et al., 2023): A two-stage modeling approach under SIM architecture, aligning computation methods of GSU and ESU to enhance consistency between them. User history length used is 15,000.
  • TWIN V2(Si et al., 2024): This approach uses hierarchical clustering to reduce scale of ultra-long user historical lengths, followed by modeling using TWIN method. User history length used is 100,000.
  • HSTU*(Zhai et al., 2024): This method employs an HSTU-style multi-layer self-masked attention mechanism to model user history (time complexity $\mathcal{O}(n^{2}*d)$). Due to computational limitations, each historical item only uses a few features: item ID, author ID, tag, and user feedback, with 2000 history length and 4 depth. Its FLOPs can be seen as the un-cache version of MARM.
@@
  • Ensemble experiment: After a modification is verified to be effective, such method will be fused with later modifications.
3.3.2. Individual Experiment Performance Comparisons
@@
    +TWIN $0.8260\pm 0.00005$ $0.7156\pm 0.00006$ 94.6M
      +TWIN V2 $0.8268\pm 0.00023$ $0.7167\pm 0.00011$ 154.6M
        +HSTU* $0.8268\pm 0.00017$ $0.7168\pm 0.00021$ 1.04B
@@
Table 3. Online A/B testing results of Short-Video services
@@
3.3.3. Ensemble Experiment Offline Performance
@@
This section aims to answer the following question: could MARM module further improve a model that already incorporates various long-term and short-term user behavior modeling? Therefore, we conducted another set of ensemble experiments and the results are shown in Table 2. Specifically, we sequentially added DIN, SIM Soft, TWIN, TWIN V2, and MARM modules to the baseline, observing improvements brought by them. From Table 2, after adding a series of long-term and short-term modeling modules, adding a simplified HSTU-style module provided no significant improvements, possibly because the scale of the features we used, attention depth, or historical length did not reach a critical mass. Besides, another reason might be that our MARM could save more interaction sample knowledge. While a pure item sequence has only a few tens of attributes, MARM’s cached results are always thousand-dimensional compressed target item query information, which achieves potential sample communication within MARM. In addition, directly adding an HSTU-style module significantly increased the computational burden on our model. In our scenario, when we add MARM to a system that includes several long-term and short-term sequence modeling models, there is still a significant improvement in accuracy. MARM is a very practical module that can be added to most recommendation models at a very controllable cost.
3.3.4. GSU searched Top-K Overlapped Rate of MARM
@@
First, we aim to understand how the attention depth of MARM leads to sustained improvements. In the framework of MARM with SIM, each layer of MARM Block models GSU and ESU phases like TWIN. The GSU returns top-k user histories most relevant to target item. As shown in Figure 6, we analyzed the overlap of the histories returned by GSU of four-layer MARM Block and found that the proportion of independence for each layer is relatively high, exceeding 50%. Moreover, the direct overlap between any two layers is less than 20%. This indicates that each layer of MARM focuses on different historical content, forming a high-level expression of interests. In addition, we elaborate on the innovation and effectiveness analysis of the cache strategy used in MARM in the Sec  4.
3.3.5. MARM Cost Discussion
@@
As an equivalent form of an un-cached MARM, HSTU follows a first sample-aggregation then learning paradigm to alleviate calculation pressure, its overall computational complexity for a user with a lifelong historical behavior is about $O(L*n^{2}*d*(\frac{N}{r}))$, where N denotes total interaction amount of a user, and $\textit{r}<<n$ is aggregation level of user samples, meaning that every r user samples are aggregated together for training. In practice, a larger aggregation r typically reduces the modeling complexity but can lead to delays in user feedback entering the model training, and easily results in a decline in online metrics, creating a trade-off between cost and effectiveness.
@@
Compared to above, HSTU applies a high-FLOPs masked-self-attention; our MARM is a more low-resource solution, needing no aggregation of training samples for user modeling. For overall learning process of a user, the computational complexity of MARM is about O(Lnd)O(L*n*d). As a replacement, MARM does require additional storage, which we previously referred to as cache size $\mathcal{C}$. In our scenario, with attention depth $L=4$ and sequence length $n=6000$, the storage used by MARM is 60TB. The combined storage and increased computational overhead is approximately $1/8$ that of a direct multi-layer self-attention approach.
For the deployed MARM version with parameters $L=4,n=6000,d=128$, serving 30M DAU, MARM uses only 100 A10 GPUs ( ¥3.5M/year ) + 60TB Storage ( ¥1.2M/year ); these resources encompass all aspects of training and inference. The sota HSTU would require at least 10 times computing resources of MARM, i.e., 1000 A10 GPUs( ¥35M/year ). The storage cost is roughly 1/3 of the additional computation cost, which is why MARM’s approach of exchanging storage for computation is so cost-effective.
Figure 7. The online ranking time cost of the deployed version of MARM is approximately 15ms, with the embedding pulling taking about 5ms and the forward computation for inference taking about 10ms.
In addition, we show in Figure 7 additional time cost that MARM brings to inference service. Compared to performance improvement provided by MARM, the time cost increase (+15ms) is a worthwhile trade-off: on par with TWIN (+15ms) and much better than HSTU (+40ms).
3.3.6. Online Results
@@
To quantify the contribution of the MARM model to real-world RecSys, we implemented it across the Retrieval, Cascading, and Ranking stages, validating its effectiveness through our online A/B testing system. The model’s performance was assessed based on key metrics, including core playtime metrics and interaction metrics such as average watch time per use and the number of likes. Table 3 presents the online results of MARM across various recommendation stages.
@@
Our app’s retention metrics are positive, indicating that the launch of MARM has resulted in a better user experience. Notably, during multi-stage, MARM achieved a significant increase in average user app usage time of +2.079%, highlighting its crucial role in enhancing user watch time. While there were slight negative impacts on Comment/Forward/Follow, significant gains were achieved in user interactions through Like. Since the primary goal of the recommendation system is to maximize average watch time, the online results are still within a reasonable range of substitution. MARM has already achieved significant business benefits in real-world recommendation systems and has been deployed to serve tens of millions of users for over a year. Moreover, through continuous iterations, MARM has brought over a 5% increase in user app usage time to our business.
4. Cache Strategy Discussion
4.0.1. Innovation Description
Actually, the NLP KV cache technique is typically applied in the inference stage of autoregressive LLM to reduce the time complexity of generating the next token. Nevertheless, our MARM cache technique is used in both the training stage and the inference stage. Overall, we have the following insightful and crucial innovations:
  • Cache Generation: Different from NLP KV cache technique which generates cache results while predicting the next token in model inference process, our MARM generates the cache results in training process, which aims to save all user-item behavior pattern information. In our MARM inference, which is shared with training-produced cache results, we focus on utilizing the saved cache results for better prediction.
  • Cache Life-Cycle: In particular, the NLP KV Cache for current sentence generation only needs temporary life-cycle storage in GPU memory. Consequently, once the sentence is generated completely, its corresponding cached data is no longer necessary, and the cache results will be deleted. Unlike language sentences that have endpoints, in streaming recommendation systems, user-item behaviors are generated sequentially, and our users do not have an ‘endpoint’ to finish. Thus, our MARM Cache should have a long-term life-cycle to store user interests, to ensure that our model can access those cached results at any time to provide high-quality recommendation accuracy.
  • Cache Scope: In NLP KV Cache, it focuses on caching all tokens’ transformed Key-results and Value-results, waiting for the incoming Query to aggregate them. In our MARM, we do not cache the transformed Key-results and Value-results, but cache the final calculated Query outputs, which could reduce the overall storage requirement, conserving resources efficiently.
  • Complexity and Resoruces: In the fields of NLP and recommendation, similar generative models like HSTU exhibit a computational complexity of $\mathcal{O}(n^{2})$ for both training and inference. Although these models utilize techniques such as kv cache during inference, they still need to compute the self-attention for the entire sequence at least once. In contrast, MARM achieves a computational complexity of $\mathcal{O}(n)$ for both training and inference because it caches the computation results for each layer in the sequence. However, MARM does incur the additional burden of requiring an external storage resource.
4.0.2. Effectiveness Analysis
Compared to computation methods using complete transformer structures like HSTU, the cache-based method of MARM is essentially equivalent in terms of computation, with the main differences being the following two points:
  • Since the key/value pairs of each layer in MARM are read from the cache, this means that the key/value pairs between two consecutive attention layers in MARM are not connected as nodes in the computation graph, only the query is connected. This implies that during backpropagation, the gradient does not pass between the key/value pairs; it only passes from the query part of the later layer to the key/value part of the earlier layer.
  • The cache in MARM is updated only once, specifically during the computation for a particular sample when it is trained and written to the cache. The portions written to the cache will be frozen, while certain model parameters retained in the computation graph, such as the Q, K and V mapping matrices and the FFN parameters for each layer, will continue to be updated.
We would like to explain why the two differences mentioned above do not significantly impair the performance of the MARM architecture. In Streaming RecSys, items that have already been exposed to a specific user will not be exposed again. Therefore, recommendation scenarios typically train using one epoch of streaming data, and repeatedly training on the same sample usually does not lead to performance improvement and may even harm the recommendation effectiveness. As a result, recommendation models remain in a state of high generalization after convergence, allowing them to handle newly emerging items and accurately model item characteristics.
Regarding the first difference, if the model already exhibits strong generalization and can accurately model the user-item pairs, it can be stored and used directly without the need for extensive adjustments during subsequent training. Instead, it is sufficient to adjust only certain parameters retained in the computation graph.
For the second difference, our analysis of the lower-level model parameters in recommendation models reveals that their changes are relatively slow, whereas the parameters closer to the logits change more rapidly. Therefore, freezing these parameters will not quickly lead to drift. Moreover, in the streaming recommendation scenario, each item often has its own lifecycle, and using the computed results from frozen parameters during its lifecycle may not be significantly worse than using the latest model parameters. This is why the cache-based MARM method still shows excellent performance and scalability in streaming recommendation scenarios.